OpenStudy (anonymous):

What fraction is the same as 62/8? Not Here 8 1/4 7 3/8 7 1/8 7 5/8

OpenStudy (anonymous):

7 3/8

sam (.sam.):

\[\frac{62}{8}=\frac{31}{4}=7\frac{3}{4}\]

sam (.sam.):

Not Here

OpenStudy (anonymous):

Agreed sam. 73/4 or 7 6/8

OpenStudy (anonymous):

not here the answer 7 6/8